OpenStudy (anonymous):

Write an equation in standard form with integer coefficients for the line with slope 18/17 going through the point (-6,-3) The equation of the line is?

OpenStudy (anonymous):

in point slope form, y + 3 = (18/17)(x + 6) so, clearing the fraction, multiply both sides by 17: 17y + 17*3 = 18(x + 6) 17y + 51 = 18x + 18*6 17y + 51 = 18x + 108 0 = 18x -17y +108 -51 0 = 18x -17y + 57... general form -57 = 18x -17y ... standard form